# Optimal Convergence Rate for Mirror Descent Methods with special Time-Varying Step Sizes Rules

### v5

@@ 5. Mirror descent method for problems with functional constraints
$\leq\frac{1}{|I|}\left(\frac{M\theta\sqrt{N}}{\sqrt{2\sigma}}+\frac{M}{\sqrt{2\sigma}}\sum_{k\in I}\frac{1}{\sqrt{k}}+\frac{M}{\sqrt{2\sigma}}\sum_{k\in J}\frac{1}{\sqrt{k}}\right)$
$=\frac{M}{|I|\sqrt{2\sigma}}\left(\theta\sqrt{N}+\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)\leq\frac{M}{|I|\sqrt{2\sigma}}\left(\theta\sqrt{N}+2\sqrt{N}\right)$
$=\frac{M\sqrt{N}(2+\theta)}{|I|\sqrt{2\sigma}}.$
Now, by setting $\frac{M\sqrt{N}(2+\theta)}{|I|\sqrt{2\sigma}}\leq\varepsilon$ and since $|I|\leq N$, we get
$\frac{M(2+\theta)}{\sqrt{N}\sqrt{2\sigma}}\leq\varepsilon\quad\Longrightarrow\quad N\geq\frac{M^{2}(2+\theta)^{2}}{2\sigma\varepsilon^{2}}.$
@@
Let $f$ and $g$ are Lipschitz convex functions with constants $M_{f}>0$ and $M_{g}>0$, respectively. Assume that $\|\nabla f(x)\|_{*}\leq M_{f}$ and $\|\nabla g(x)\|_{*}\leq M_{g}$, for all $x\in Q$, and $V_{\psi}(x^{*},x^{1})\leq\theta,$ for some $\theta>0$, and let $\varepsilon>0,M:=\max\{M_{f},M_{g}\}$. Then after
$N=\left\lceil\frac{M^{2}(2+\theta)^{2}}{2\sigma\varepsilon^{2}}\right\rceil=O\left(\frac{1}{\varepsilon^{2}}\right),$
@@
$f(\tilde{x})-f(x^{*})$ $\leq\frac{M}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\sqrt{k}}\cdot\left(\theta+\sum_{k=1}^{N}\frac{1}{k}\right)$
$\leq\frac{M}{\sqrt{2\sigma}}\cdot\frac{1}{2\sqrt{|I|+1}-2}\cdot\left(\theta+1+\log(N)\right)$
$\leq\frac{M\left(\theta+1+\log(N)\right)}{\sqrt{\sigma}\sqrt{|I|}}.$
@@
For solving problem (5.1) (or its equivalence (5.2)), when we have a very big number of functional constraints, we can modify Algorithm 3 in a similar way as in [31]. When we have a non-productive step $k$, i.e., it holds that $g(x^{k})>\varepsilon$, then instead of calculating a (sub)gradient of the functional constraint with max-type $g(x)=\max_{1\leq i\leq p}\{g_{i}(x)\}$, we calculate a (sub)gradient of one functional $g_{i}$, for which we have $g_{i}(x^{k})>\varepsilon$. This idea of the modification of Algorithms 3 allows saving the running time of the algorithm due to consideration of not all functional constraints (when their number is big) on non-productive steps.
@@ 6. Numerical experiments
We compare the performance of the studied algorithms with the projected subgradient method using different famous step size rules that are listed in Table 1. Therefore, in our experiments, we take the standard Euclidean prox-structure, namely $\psi(x)=\frac{1}{2}\|x\|_{2}^{2}$ which is $1$-strongly functions (i.e., $\sigma=1$) and the corresponding Bregman divergence is $V_{\psi}(x,y)=\frac{1}{2}\|x-y\|_{2}^{2}$. In all experiments, we take the set $Q$ as a unit ball in $\mathbb{R}^{n}$ with the center at $0\in\mathbb{R}^{n}$. All compared methods start from the same initial point $x^{1}=\left(\frac{1}{\sqrt{n}},\ldots,\frac{1}{\sqrt{n}}\right)\in Q\subset\mathbb{R}^{n}$. In the AdaGrad algorithm, we take $\alpha=10^{-8}$, and there is an assumption that $\|x^{1}-x^{*}\|_{2}^{2}\leq 2\theta_{0}^{2}$, thus for the taken feasible set in our experiments, we can take $\theta_{0}=1/\sqrt{2}$.
@@
new non-adaptive this paper (3.7) $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}}$,
new adaptive this paper (3.7) $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}}$.
@@ 6.4. Maximum of a finite collection of linear functions
(6.4) $\min_{Q}f(x):=\max\left\{f_{i}(x)=\langle a_{i},x\rangle+b_{i},i=1,2,\ldots,T\right\},$

### v6

@@ 5. Mirror descent method for problems with functional constraints
$\leq\frac{1}{|I|}\left(\frac{M\theta_{1}\sqrt{N}}{\sqrt{2\sigma}}+\frac{M}{\sqrt{2\sigma}}\sum_{k\in I}\frac{1}{\sqrt{k}}+\frac{M}{\sqrt{2\sigma}}\sum_{k\in J}\frac{1}{\sqrt{k}}\right)$
$=\frac{M}{|I|\sqrt{2\sigma}}\left(\theta_{1}\sqrt{N}+\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)\leq\frac{M}{|I|\sqrt{2\sigma}}\left(\theta_{1}\sqrt{N}+2\sqrt{N}\right)$
$=\frac{M\sqrt{N}(2+\theta_{1})}{|I|\sqrt{2\sigma}}.$
Now, by setting $\frac{M\sqrt{N}(2+\theta_{1})}{|I|\sqrt{2\sigma}}\leq\varepsilon$ and since $|I|\leq N$, we get
$\frac{M(2+\theta_{1})}{\sqrt{N}\sqrt{2\sigma}}\leq\varepsilon\quad\Longrightarrow\quad N\geq\frac{M^{2}(2+\theta_{1})^{2}}{2\sigma\varepsilon^{2}}.$
@@
Let $f$ and $g$ are Lipschitz convex functions with constants $M_{f}>0$ and $M_{g}>0$, respectively. Assume that $\|\nabla f(x)\|_{*}\leq M_{f}$ and $\|\nabla g(x)\|_{*}\leq M_{g}$, for all $x\in Q$, and $V_{\psi}(x^{*},x^{1})\leq\theta_{1},$ for some $\theta_{1}>0$, and let $\varepsilon>0,M:=\max\{M_{f},M_{g}\}$. Then after
$N=\left\lceil\frac{M^{2}(2+\theta_{1})^{2}}{2\sigma\varepsilon^{2}}\right\rceil=O\left(\frac{1}{\varepsilon^{2}}\right),$
@@
$f(\tilde{x})-f(x^{*})$ $\leq\frac{M}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\sqrt{k}}\cdot\left(\theta_{1}+\sum_{k=1}^{N}\frac{1}{k}\right)$
$\leq\frac{M}{\sqrt{2\sigma}}\cdot\frac{1}{2\sqrt{|I|+1}-2}\cdot\left(\theta_{1}+1+\log(N)\right)$
$\leq\frac{M\left(\theta_{1}+1+\log(N)\right)}{\sqrt{\sigma}\sqrt{|I|}}.$
@@
5.1. Modification of mirror descent method for problems with many functional constraints
For solving problem (5.1), when we have several functional constraints, we can modify Algorithm 3 in a similar way as in [31]. When we have a non-productive step $k$, i.e., it holds that $g(x^{k})>\varepsilon$, then instead of calculating a (sub)gradient of the functional constraint with max-type $g(x)=\max_{1\leq i\leq p}\{g_{i}(x)\}$, we calculate a (sub)gradient of one functional $g_{i}$, for which we have $g_{i}(x^{k})>\varepsilon$. This idea of the proposed modification allows for saving the running time of the algorithm due to consideration of not all functional constraints (when their number is big) on non-productive steps. The proposed modification is listed as Algorithm 4 below.
Algorithm 4 Modification of mirror descent method for problems with many functional constraints.
0:  $\varepsilon>0$, initial point $x^{1}\in Q$, $M=\max\{M_{f},M_{g}\}$, $\theta_{1}>0$ such that $V_{\psi}(x^{*},x^{1})\leq\theta_{1}$, $m\geq-1$, $\sigma>0$.
1:  $I\longrightarrow\emptyset,J\longrightarrow\emptyset.$
2:  Set $k=0$
3:  repeat
4:     if $g_{i}(x^{k})\leq\varepsilon,\;\forall i=1,2,\ldots,p$ then
5:        $L_{k}=\|\nabla f(x^{k})\|_{*},$
6:        $\gamma_{k}^{f}=\frac{\sqrt{2\sigma}}{L_{k}\sqrt{k}},$
7:        $x^{k+1}=\arg\min_{x\in Q}\left\{\langle x,\nabla f(x^{k})\rangle+\frac{1}{\gamma_{k}^{f}}V_{\psi}(x,x^{k})\right\},$
8:        $k\longrightarrow I$   ”productive step”
9:     else
10:        (i.e., $\exists q=q(k)\in\{1,2,\ldots,p\}$, s.t., $g_{q(k)}>\varepsilon$)
11:        $L_{k}=\|\nabla g_{q(k)}(x^{k})\|_{*},$
12:        $\gamma_{k}^{g_{q(k)}}=\frac{\sqrt{2\sigma}}{L_{k}\sqrt{k}},$
13:        $x^{k+1}=\arg\min_{x\in Q}\left\{\langle x,\nabla g_{q(k)}(x^{k})\rangle+\frac{1}{\gamma_{k}^{g_{q(k)}}}V_{\psi}(x,x^{k})\right\},$
14:        $k\longrightarrow J$   ”non-productive step”
15:     end if
16:  until
$\varepsilon\sum_{i=1}^{k}\left(\frac{L_{i}\sqrt{i}}{\sqrt{2\sigma}}\right)^{m}\geq\theta_{1}\left(\frac{M\sqrt{k}}{\sqrt{2\sigma}}\right)^{m+1}$ $+\frac{1}{\left(\sqrt{2\sigma}\right)^{m+1}}\Bigg{(}\sum_{i\in I}\left(\sqrt{i}\right)^{m-1}\|\nabla f(x^{i})\|_{*}^{m+1}$
$\qquad\qquad+\sum_{j\in J}\left(\sqrt{j}\right)^{m-1}\|\nabla g_{q(j)}(x^{j})\|_{*}^{m+1}\Bigg{)}.$
16:  $\hat{x}=\frac{1}{\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}}\sum\limits_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}x^{i}.$
For Algorithm 4, we have the following result.
Theorem 5.7.
Let $f$ and $g_{i}\,(i=1,2,\ldots,p)$ are Lipschitz convex functions with constants $M_{f}>0$ and $M_{g_{i}}>0\,(i=1,2,\ldots,p)$, respectively. Assume that, for all $x\in Q$, we have $\|\nabla f(x)\|_{*}\leq M_{f}$ and $\|\nabla g_{i}(x)\|_{*}\leq M_{g}\,\forall i=1,2,\ldots,p$, where $M_{g}=\max_{1\leq i\leq p}\{M_{g_{i}}\}$, and $V_{\psi}(x^{*},x^{1})\leq\theta_{1}$, for some $\theta_{1}>0$, and let $\varepsilon>0$, $M=\max\{M_{f},M_{g}\}$. Then for problem (5.1), after stopping Algorithm 4, for any fixed $m\geq-1$, it satisfies the following inequalities
(5.14) $f(\hat{x})-f(x^{*})<\varepsilon,\quad\text{and}\quad g_{i}(\hat{x})\leq\varepsilon,\,\forall i=1,\ldots,p.$
where $\hat{x}=\frac{1}{\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}}\sum\limits_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}x^{i}$, and $x^{*}$ is an optimal solution of (5.1).
Proof.
Similar to what was done in proving Theorem 5.1, we find that for all $i\in I$,
(5.15) $\frac{f(x^{i})-f(x^{*})}{\left(\gamma_{i}^{f}\right)^{m}}\leq\frac{V_{\psi}(x^{*},x^{i})-V_{\psi}(x^{*},x^{i+1})}{\left(\gamma_{i}^{f}\right)^{m+1}}+\frac{\|\nabla f(x^{i})\|_{*}^{2}}{2\sigma\left(\gamma_{i}^{f}\right)^{m-1}},$
and for all $j\in J$, there is $q(j)\in\{1,2,\ldots,p\}$, such that
(5.16) $\frac{g(x^{j})-g(x^{*})}{\left(\gamma_{j}^{g_{q(j)}}\right)^{m}}\leq\frac{V_{\psi}(x^{*},x^{j})-V_{\psi}(x^{*},x^{j+1})}{\left(\gamma_{j}^{g_{q(j)}}\right)^{m+1}}+\frac{\|\nabla g_{q(j)}(x^{j})\|_{*}^{2}}{2\sigma\left(\gamma_{j}^{g_{q(j)}}\right)^{m-1}}.$
By taking the summation, in each side of (5.15) and (5.16), over productive and non-productive steps, with $\gamma_{i}=\gamma_{i}^{f}$ if $i\in I$ and $\gamma_{j}=\gamma_{j}^{g_{q(j)}}$ if $i\in J$, we get
$\quad\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}\left(f(x^{i})-f(x^{*})\right)+\sum_{j\in J}\left(\gamma_{j}^{g_{q(j)}}\right)^{-m}\left(g_{q(j)}(x^{j})-g_{q(j)}(x^{*})\right)$
$\leq\sum_{i=1}^{k}\frac{1}{\gamma_{i}^{m+1}}\left(V_{\psi}(x^{*},x^{i})-V_{\psi}(x^{*},x^{i+1})\right)+\frac{1}{2\sigma}\sum_{i\in I}\frac{\|\nabla f(x^{i})\|_{*}^{2}}{\left(\gamma_{i}^{f}\right)^{m-1}}$
$\;\;\;\;+\frac{1}{2\sigma}\sum_{j\in J}\frac{\|\nabla g_{q(j)}(x^{j})\|_{*}^{2}}{\left(\gamma_{j}^{g_{q(j)}}\right)^{m-1}}.$
Since for any $j\in J$, it holds
$g_{q(j)}(x^{j})-g_{q(j)}(x^{*})\geq g_{q(j)}(x^{j})>\varepsilon>0,$
and since $f$ is a convex function, we get
$\quad\left(\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}\right)\left(f(\hat{x})-f(x^{*})\right)$
$<\frac{\theta_{1}}{\gamma_{k}^{m+1}}+\frac{1}{2\sigma}\sum_{i\in I}\frac{\|\nabla f(x^{i})\|_{*}^{2}}{\left(\gamma_{i}^{f}\right)^{m-1}}+\frac{1}{2\sigma}\sum_{j\in J}\frac{\|\nabla g_{q(j)}(x^{j})\|_{*}^{2}}{\left(\gamma_{j}^{g_{q(j)}}\right)^{m-1}}-\varepsilon\sum_{j\in J}\left(\gamma_{j}^{g_{q(j)}}\right)^{-m}$
$=\frac{\theta_{1}}{\gamma_{k}^{m+1}}+\frac{1}{2\sigma}\sum_{i\in I}\frac{\|\nabla f(x^{i})\|_{*}^{2}}{\left(\gamma_{i}^{f}\right)^{m-1}}+\frac{1}{2\sigma}\sum_{j\in J}\frac{\|\nabla g_{q(j)}(x^{j})\|_{*}^{2}}{\left(\gamma_{j}^{g_{q(j)}}\right)^{m-1}}-\varepsilon\sum_{i=1}^{k}(\gamma_{i})^{-m}$
$\quad+\varepsilon\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}.$
But $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}}$ or $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla g_{q(k)}(x^{k})\|_{*}\sqrt{k}}$ and in both cases we have $(\gamma_{k})^{m+1}\geq\left(\frac{\sqrt{2\sigma}}{M\sqrt{k}}\right)^{m+1}$. Thus, we have
$\quad\left(\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}\right)\left(f(\hat{x})-f(x^{*})\right)$
$<\varepsilon\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}+\theta_{1}\left(\frac{M\sqrt{k}}{\sqrt{2\sigma}}\right)^{m+1}+\frac{1}{2\sigma}\sum_{i\in I}\|\nabla f(x^{i})\|_{*}^{2}\left(\frac{\|\nabla f(x^{i})\|_{*}\sqrt{i}}{\sqrt{2\sigma}}\right)^{m-1}$
$\quad+\frac{1}{2\sigma}\sum_{j\in J}\|\nabla g_{q(j)}(x^{j})\|_{*}^{2}\left(\frac{\|\nabla g_{q(j)}(x^{j})\|_{*}\sqrt{j}}{\sqrt{2\sigma}}\right)^{m-1}-\varepsilon\sum_{i=1}^{k}\left(\frac{L_{i}\sqrt{i}}{\sqrt{2\sigma}}\right)^{m}$
$\quad=\varepsilon\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}-\Bigg{(}\varepsilon\sum_{i=1}^{k}\left(\frac{L_{i}\sqrt{i}}{\sqrt{2\sigma}}\right)^{m}-\theta_{1}\left(\frac{M\sqrt{k}}{\sqrt{2\sigma}}\right)^{m+1}$
$\qquad\qquad\qquad\qquad\qquad-\frac{1}{(\sqrt{2\sigma})^{m+1}}\sum_{i\in I}(\sqrt{i})^{m-1}\|\nabla f(x^{i})\|_{*}^{m+1}$
$\qquad\qquad\qquad\qquad\qquad-\frac{1}{(\sqrt{2\sigma})^{m+1}}\sum_{j\in J}(\sqrt{j})^{m-1}\|\nabla g_{q(j)}(x^{j})\|_{*}^{m+1}\Bigg{)}.$
Therefore, from stopping criterion of Algorithm 4, we get
$\left(\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}\right)\left(f(\hat{x})-f(x^{*})\right)<\varepsilon\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}.$
Thus, by dividing both sides of the previous inequality by $\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}$, we get the desired inequality $f(\hat{x})-f(x^{*})<\varepsilon$.
For all $i\in I$ it holds that $g_{s}(x^{i})\leq\varepsilon\,\forall s\in\{1,2,\ldots,p\}$, and since $g_{s}\,(\forall s\in\{1,2,\ldots,p\})$ are convex functions, then we have
$g_{s}(\hat{x})\leq\frac{1}{\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}}\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}g_{s}(x^{i})\leq\varepsilon,\;\forall s\in\{1,2,\ldots,p\}.$
@@ 5.1. Modification of mirror descent method for problems with many functional constraints
For any $m\geq 1$, let us assume that Algorithm 4 works $N$ iterations. Then we have
$\qquad f(\hat{x})-f(x^{*})$
$<\frac{\left(\sqrt{2\sigma}\right)^{m}}{M^{m}\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\Bigg{(}\theta_{1}\left(\frac{M\sqrt{N}}{\sqrt{2\sigma}}\right)^{m+1}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{M^{m+1}\left(\sqrt{k}\right)^{m-1}}{\left(\sqrt{2\sigma}\right)^{m-1}}\Bigg{)}$
$=\frac{M}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\cdot\left(\theta_{1}\left(\sqrt{N}\right)^{m+1}+\sum_{k=1}^{N}\left(\sqrt{k}\right)^{m-1}\right)$
$\leq\frac{M}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\cdot\left(\theta_{1}\left(\sqrt{N}\right)^{m+1}+N\left(\sqrt{N}\right)^{m-1}\right)$
$=\frac{M(1+\theta_{1})\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}.$
Now, by setting $\frac{M(1+\theta_{1})\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\leq\varepsilon$ and since $|I|\leq N$, we get
$\frac{M(1+\theta_{1})\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}N\left(\sqrt{N}\right)^{m}}\leq\frac{M(1+\theta_{1})\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\leq\varepsilon.$
Thus,
$\frac{M(1+\theta_{1})}{\sqrt{N}\sqrt{2\sigma}}\leq\varepsilon\quad\Longrightarrow\quad N\geq\frac{M^{2}(1+\theta_{1})^{2}}{2\sigma\varepsilon^{2}}.$
Therefore, we find that after
$N=\left\lceil\frac{M^{2}(1+\theta_{1})^{2}}{2\sigma\varepsilon^{2}}\right\rceil=O\left(\frac{1}{\varepsilon^{2}}\right)$
iterations of Algorithm 4, for any fixed $m\geq 1$, it satisfy $f(\hat{x})-f(x^{*})<\varepsilon$ and $g_{i}(\hat{x})\leq\varepsilon\,\forall i\in\{1,2,\ldots,p\}$.
@@ 6. Numerical experiments
First, we mention that for the following adaptive step size scheme
(6.1) $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}}\quad k=1,\ldots N,$
there is no guarantee that the sequence $\{\gamma_{k}\}_{k\geq 1}$ is non-increasing. As a result, this step size does not satisfy one of the conditions of Theorem 3.1. However, Algorithm 1 performs well in practice, and the convergence rate remains the same as the step size defined in (3.7). Therefore, we need to consider restructuring the scheme of step sizes so that it becomes adaptive and allows us to obtain the same results as previously obtained for the non-adaptive ones.
We compare the performance of the studied algorithms with the projected subgradient method using different famous step size rules that are listed in Table 1. Therefore, in our experiments, we take the standard Euclidean prox-structure, namely $\psi(x)=\frac{1}{2}\|x\|_{2}^{2}$ which is $1$-strongly functions (i.e., $\sigma=1$) and the corresponding Bregman divergence is $V_{\psi}(x,y)=\frac{1}{2}\|x-y\|_{2}^{2}$. In all experiments, we take the set $Q$ as a unit ball in $\mathbb{R}^{n}$ with the center at $0\in\mathbb{R}^{n}$. All compared methods start from the same initial point $x^{1}=\left(\frac{1}{\sqrt{n}},\ldots,\frac{1}{\sqrt{n}}\right)\in Q\subset\mathbb{R}^{n}$. In the AdaGrad algorithm, we take $\alpha=10^{-8}$, and there is an assumption that $\|x^{1}-x^{*}\|_{2}^{2}\leq 2\theta_{0}^{2}$, thus for the taken feasible set in our experiments, we can take $\theta_{0}=\sqrt{2}$.
@@
non-adaptive time-varying scheme (3.7) $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}}$,
adaptive time-varying scheme (6.1) $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}}$.
@@ 6.4. Maximum of a finite collection of linear functions
(6.5) $\min_{x\in Q}\left[f(x):=\max\left\{f_{i}(x)=\langle a_{i},x\rangle+b_{i},i=1,2,\ldots,T\right\}\right],$